OpenStudy (anonymous):

A tool box has a volume of x3 + 8x2 + 11x – 20 cm3 and the height is x + 5 cm. Find the polynomial that would represent the area of the bottom of the tool box? Explain your reasoning.

OpenStudy (anonymous):

\(v=x^3+8x^2+11x-20\) \(h=x+5\) volume = length x width x height area = length x width You are looking for the area, so substitute: volume = area x height and solve for area.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your welcome. Let me know if you want to check your answer.